OpenStudy (anonymous):

can someone teach me how to draw a circle in Mathematica Wolfram. I have the the center and the radius.

OpenStudy (anonymous):

Mathematica or wolfram alpha?

OpenStudy (anonymous):

The normal one. I have it installed at home

OpenStudy (anonymous):

is not the same btw?

OpenStudy (anonymous):

Nah, wolfram alpha is the website. Mathematica is the code based program.

OpenStudy (anonymous):

What is the circle?

OpenStudy (anonymous):

Wolfram 8 installed at home

OpenStudy (anonymous):

center (-4,-2) f=sqrt(5)

OpenStudy (anonymous):

Is f the focal or do you mean radius?

OpenStudy (anonymous):

sorry i meant to write r=sqrt(5)

OpenStudy (anonymous):

Okay, just making sure. What I would do is do this: \[(x+4)^2+(y+2)^2=5\]

OpenStudy (anonymous):

\[(y+2)^2=5-(x+4)^2\] \[y+2=\pm \sqrt{5-(x+4)^2}\] \[y=\sqrt{5-(x+4)^2}-2\] \[y=-\sqrt{5-(x+4)^2}-2\] Plot the last two. One is the top half the other is the bottom

OpenStudy (anonymous):

I did it with the following syntax: Graphics[Circle[{-4, -2}, Sqrt[5]]] However I can only see the circle with not axis. Can someone help me with this?

OpenStudy (anonymous):

Oh, mine is a little far fetched then haha.

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

Already fixed that

OpenStudy (anonymous):

A sample plot procedure written in Mathematica 8 Home Edition is attached. There is a misspelling in the comments here and there.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Glad that someone besides myself is using Mathematica on this site.

OpenStudy (anonymous):

Attached is a screen capture of my 8Gb IMac setup that I use for solving openstudy.com math problems.